k 0 1 (c) Consider the matrix 0 k 2 -2 k 3 i. Compute the determinant. ii. For what value(s) of k does A- exist? iii. For whaNeed this asap
Thanks
Answer #1

bet a= K 2 .2 K 3 (i) det () = *(3*--2K) +1(2x) = x+ 2K = *(x+2) (ii) 94 K+0, K7-2. then. det (a) #0 and hence AT! exista (ii